Tatoo Incorporated reported a net capital loss of $13,500 in 2020. The company had a net capital gain of $4,800 in 2018 and $3,500 in 2017. In 2019, although the company suffered a net operating loss, it had net capital gains of $1,500. What is the amount of Tatoo's capital loss carryover to 2021 remaining after it applies the carryback?

a) $3,700.

b) $5,200.

c) $8,700.

d) $13,500.
Answer #1
Share of Retail store in San Franciso -20,825.00 =-29750*70%
Share of remaining profits      24,661.25 =(131650+20825-(76900*70%))*25%
Income will be allocated to Reagan in total $ 3,836.25 =24661.25-20825
Correct answer is option C .)